OpenStudy (anonymous):

A rectangular prism has length of 4 feet and a width of 7 feet. If the surface area of the prism is 188 square feet, what is its height? 6 ft 5 ft 4 ft 3 ft

OpenStudy (anonymous):

2*28 + 2*4*h + 2*7*h = 188 56 + 8h +14h = 188 22h = 188-56 = 132 h = 6 height = 6 feet

OpenStudy (anonymous):

Using the surface area formula for a rectangula prism A=2(lw+lh+hw) Substitute the known values 188=2(4h+4*7+7h) grouping similar terms 188=2(11h+28) 188=22h+56 all thats left is to Solve for h